Consider the following functional relations between x and y.

Y = (2x2 – 6x-20)2

a) at what values of x will the function have zero slope?

b) identify whether those zero slope points are maximum or minimum values of the function.


1
Expert's answer
2021-12-15T11:34:50-0500

"Y = (2x^2 \u2013 6x-20)^2"

The slope is;

"\frac{\Delta Y}{\Delta x}=2(2x^2-6x-20)(4x-6)"

Putting "\frac{\Delta Y}{\Delta x}=0" ;

Either 2x2-6x-20=0 or 4x-6=0

For 2x2-6x+20=0;


"x=\frac{6\pm\sqrt{36+160}}{2×2}"

"x=5,-2"

For 4x-6;

"x=\frac{6}{4}=\frac{3}{2}"

"\therefore" for "x=-2,5,\frac{3}{2}" Slope is zero


b)For x=2. Y=0 (min)

For x=5. Y=0 (min)

For x="\frac{-3}{2}" Y=42.25 ( max)